OpenStudy (anonymous):

p=l^2r find p if l=12 amperes r-2 ohms a=2w^2 + 4hw find a if w=7cm h=10cm

OpenStudy (anonymous):

p= 12^4 a= 14^2 + 74

OpenStudy (ash2326):

p-I^2 R I=12 I^2=144 R=2 P=144*2=288 watts a=2w^2+4hw w=7 cm h=10 cm a=2*49+4*7*10 a=98+280 a=378 cm^2